# Simulated Annealing for Quadratic and Higher-Order Unconstrained Integer Optimization

### v1

@@ 4 Results
In this section, we present numerical results for SA with the transition probabilities described in Sect. 3.3. Our proposed algorithms are implemented in the open-source optimization framework OpenJij[openjij], and are freely available for use. All numerical experiments presented below were performed using OpenJij on a machine equipped with an Apple M1 Max processor and 64 GB of memory. In addition, we used the libraries JijModeling[cite_jijmodeling] and OMMX[cite_ommx_1, cite_ommx_2] to implement the QUIO and HUIO problems. Both libraries provide Python front ends with Rust back ends, enabling fast construction of problem instances.
@@
Let us summarize the basic properties of these models. The fully connected model $E^{\text{FC}}_{p,u}(\bm{z})$ consists of fully connected quadratic terms and on-site terms of degree $p$. Its ground-state energy is N(N+1)/2-N(N+1)/2. The multilinear model $E^{\text{ML}}_{p,u}(\bm{z})$ contains $p$-body multilinear interactions with periodic boundary conditions, $z_{i+N}=z_{i}$. Its ground-state energy is N-N. Both of these models have analytically known ground-state energies. In contrast, the ground-state energy of the random interaction model $E^{\text{RI}}_{p,u}(\bm{z})$ is generally difficult to determine, since it contains random $p$-body interactions whose strengths $J_{i_{1},\ldots,i_{p}}$ are sampled uniformly from [1,1][-1,1]. We set the number of interactions to 1000 for the model to be sparse. Note that this model may include on-site terms of the form $z_{i}^{k}$.
@@
4.1 Annealing time per sweep
In evaluating the performance of optimization algorithms, both solution quality and computational cost are important metrics. However, in this subsection we focus exclusively on the annealing time of SA for a fixed number of sweeps and we investigate how the annealing time, defined as the elapsed time from the start to the end of a single SA process, scales with the problem characteristics. In the following numerical experiments, SA was performed with 100 samples and 1000 sweeps, and the mean annealing time per single SA run was computed.
Let us first investigate how the variable bound range, $2u+1$, influences the annealing time, focusing on the fully connected model [Eq. (32)]. Figure 1 shows the annealing time as a function of $2u+1$, for $p=1$ and $p=2$. For $p=1$, the annealing time does not increase with $2u+1$ for any of the transition probabilities [Fig. 1(a)], because determining the update of a single variable does not depend on the variable bound range. For $p=2$, only the heat bath method requires $O(2u+1)$ time to determine the update of a single variable, because all $2u+1$ candidate values must be examined when performing the update. Consequently, as shown in Fig. 1(b), only the heat bath method exhibits annealing time that increases linearly with $2u+1$.
Figure 1: Annealing time as a function of variable bound range $2u+1$ for (a) $E^{\text{FC}}_{p=1,u}(\bm{z})$ and (b) $E^{\text{FC}}_{p=2,u}(\bm{z})$, obtained by SA with different transition probabilities. Each data point represents the mean annealing time over 100 samples, with 1000 sweeps and $N=100$ variables. Error bars indicate the standard error.
Figure 2: Annealing time as a function of $p$ for (a) $E^{\text{FC}}_{p,u=1}(\bm{z})$ and (b) $E^{\text{ML}}_{p,u=10}(\bm{z})$, obtained by SA with different transition probabilities. Each data point represents the mean annealing time over 100 samples, with 1000 sweeps and $N=100$ variables. Error bars indicate the standard error but are too small to be visible.
@@ 4.2 Energy as a function of temperature
Figure 3: Temperature $T$ dependence of the mean energy obtained by SA with different transition probabilities for (a) $E^{\text{FC}}_{p=2,u=1}(\bm{z})$, (b) $E^{\text{FC}}_{p=2,u=100}(\bm{z})$, (c) $E^{\text{RI}}_{p=4,u=1}(\bm{z})$, and (d) $E^{\text{RI}}_{p=4,u=100}(\bm{z})$. Each data point represents the mean energy at temperature $T$ during SA, averaged over 1000 samples, and the number of sweeps is 1000. The number of variables is $N=100$, and for $E^{\text{RI}}_{p,u}(\bm{z})$, the number of interactions is 1000. The black dotted line for (a) and (b) indicates the ground-state energy, 5050-5050.
@@
Similarly, Figs. 3(c) and 3(d) show the results for the random interaction model $E^{\text{RI}}_{p=4,u}(\bm{z})$. This model is a HUIO problem with fourth-order interactions and random coupling coefficients, and its characteristics therefore differ from those of the fully connected model $E^{\text{FC}}_{p=2,u}(\bm{z})$. Nevertheless, the results exhibit similar tendencies to those of $E^{\text{FC}}_{p=2,u}(\bm{z})$. The initial temperature appears appropriate, whereas the final temperature tends to be too low for this model as well.
@@
As for the differences among transition probabilities, the energy tends to decrease more readily when using the heat bath method than when using the Metropolis method [Figs. 3(b) and 3(d)], since it takes into account all possible variable updates instead of randomly selecting a single candidate. For the optimal-transition Metropolis method, the energy occasionally drops sharply when an optimal transition actually occurs, and thus its energy is generally lower than that of the other methods at the same temperature.
@@
Figure 4: Temperature $T$ dependence of the mean energy obtained by SA with different transition probabilities for $E^{\text{ML}}_{p=30,u=1}(\bm{z})$. Each data point represents the mean energy at temperature $T$ during SA, averaged over 1000 samples, and the number of sweeps is (a) 1000 and (b) 100000. The number of variables is $N=100$, and the black dotted line indicates the ground-state energy, 100-100.
@@
Next, we examine the results for the multilinear model $E^{\text{ML}}_{p=30,u=1}(\bm{z})$. For the case of 1000 sweeps [Fig. 4(a)], the energy is still decreasing even in the low-temperature region, suggesting that the final temperature is too high. In contrast, for 100000 sweeps [Fig. 4(b)], a region appears at low temperatures where the energy no longer decreases, indicating that the final temperature is sufficiently low. These results show that while the initial temperature is sufficiently high, the final temperature may be inadequate depending on the number of sweeps. In general, when the number of sweeps is small, the final temperature needs to be set lower. With fewer sweeps, variable updates are limited, so allocating more sweeps to the low-temperature region becomes important.
@@
We note that the multilinear model has a rather special nature. In each term of $E^{\text{ML}}_{p,u}(\bm{z})$, if any variable takes the value zero, the term contributes nothing to the energy. Because each term is a product of $p$ variables, the probability that a term is nonzero becomes extremely small for large $p$. For instance, when $u=1$, only $2^{p}$ configurations yield a nonzero value for a single term, whereas $3^{p}-2^{p}$ configurations yield zero. For $p=30$, this difference is enormous. As a result, there exist a vast number of degenerate states satisfying $E^{\mathrm{ML}}_{p=30,u=1}(\bm{z})=0$, and escaping from them requires a very large number of sweeps. Indeed, even with 100000 sweeps, the energy does not reach the ground-state value 100-100 [Fig. 4(b)], indicating that a much larger number of sweeps is required.
@@
In summary, the temperature-setting procedure in Sect. 3.4 is reasonable for the models studied here, although both the initial and final temperatures tend to be slightly conservative and are not necessarily optimal. In particular, the final temperature can be either too low or too high. Ideally, these temperatures should be chosen not only from the model itself but also by taking factors such as the number of sweeps into account.
@@ 4.3 Energy as a function of sweeps
In this subsection, we examine how the energy decreases as a function of the number of sweeps. Since the computational cost of one sweep depends on the choice of transition probability, the sweep–energy relationship alone does not quantify performance, but it helps reveal the characteristic behavior of each transition probability. In the following experiments, SA was performed with 1000 samples, and the mean energy for each fixed number of sweeps was computed.
We first consider the random interaction model [Eq. (34)] for $p=2$. When the bound is small ($u=1$), the three transition probability methods exhibit almost no difference in performance [Fig. 5(a)]. In contrast, when the bound is increased to $u=100$, the Metropolis method yields noticeably higher energies, particularly at small sweep regimes [Fig. 5(b)]. This behavior is consistent with the discussion in Sect. 3.3.1, which states that the Metropolis method performs poorly when the variable bounds are wide. It is also observed that the heat-bath method achieves slightly lower energies than the optimal-transition Metropolis method in this regime [Fig. 5(b)].
@@
The results for the multilinear model [Eq. (33)] show a somewhat different trend. For the extremely large bound $u=10^{6}$, the Metropolis method again performs poorly, but the heat-bath method exhibits a temporary degradation at small sweeps [Fig. 6(a)]. Fortunately, these small-sweep regimes are computed almost instantaneously, so one can practically run far more sweeps and this behavior is not a practical issue. For the case of large polynomial degree ($p=30$), the system does not reach the ground state even after $10^{6}$ sweeps for any transition probability, and in this regime the optimal-transition Metropolis method yields the lowest energies.
Figure 6: Sweep dependence of the mean energy obtained by SA with different transition probabilities for (a) $E^{\text{ML}}_{p=2,u=10^{6}}(\bm{z})$ and (b) $E^{\text{ML}}_{p=30,u=1}(\bm{z})$, Each data point represents the mean energy averaged over 1000 samples. Error bars indicate the standard error but are too small to be visible. The number of variables is $N=100$ and the black dotted line indicates the ground-state energy, 100-100.
In summary, in terms of energy per sweep, the Metropolis method shows poor performance when the variable bounds are large. For the remaining two methods, depending on the sweep regime, either the heat-bath method or the optimal-transition Metropolis method can perform slightly better, but the difference between them remains small.
@@ 4.4 Energy as a function of annealing time
In this subsection, we examine the dependence between the optimized energy and the annealing time for different transition probabilities. This analysis aims to evaluate the performance of each transition probability, where better performance corresponds to achieving a lower energy within a shorter annealing time. In addition, we compare these results with those obtained by applying SA with the Metropolis method to QUBO and HUBO formulations, which are constructed by converting the integer variables into binary ones using the log encoding:
@@
Here, $m_{k}=\lceil\log_{2}(u_{k}-l_{k}+1)\rceil$ is the number of binary variables $b_{k,i}\in\{0,1\}$ corresponding to an integer variable $z_{k}$. Although the total computational cost should include the time required for binary encoding, we exclude it here to focus on the dependence of the achieved energy on the annealing time. It should be noted, however, that for models with higher-order interactions or large variable bounds, such binary conversion can require a substantial amount of time. In the following analysis, SA was performed with 1000 samples for several fixed numbers of sweeps, and the mean energy obtained in each case was plotted against the mean annealing time.
@@
First, we examine the results for the random interaction model [Eq. (34)]. Figure 7(a) shows the results for $p=2$ and $u=1$. When this model is converted into a QUBO formulation using the log encoding, the number of variables becomes twice as large and the number of interactions becomes four times larger, which increases the computational cost per sweep. As a result, SA applied to the log-encoded QUBO formulation yields higher energies at short annealing times, indicating poorer solution quality. In contrast, little difference is observed among the other transition probabilities, because when $u=1$, each variable can take only three states (1,0-1,0, and $1$), so the influence of the choice of transition probability become less significant. As the annealing time increases, the optimized energies obtained by all methods converge to similar levels, suggesting that all methods tend to reach the global optimum when sufficient annealing time is allowed.
Next, we examine the results for larger $u=100$ [Fig. 7(b)]. In this case, the difference between the log-encoded QUBO formulation and the directly solved QUIO model becomes larger than in the case of $u=1$. This is simply because both the number of variables and the number of interaction terms become much larger than in the case of $u=1$. The heat bath method also exhibits slower energy convergence with respect to annealing time, because approximately 1% of the interactions are squared terms such as $z^{2}$, and determining the updates for these variables requires $O(2u+1)$ time in the heat bath method. The optimal-transition Metropolis method shows lower energies in the short-time region and reaches the same final energies as the other methods in the long-time region, indicating that it provides the best overall performance. This is due to the effect of the optimal transitions defined in Eq. (25).
Similar tendencies are also observed for the HUIO case. As in the $p=2$ case, a clear difference is observed between the log-encoded HUBO formulation and the directly solved HUIO model [Figs. 7(c) and 7(d)]. In general, the number of variables scales as $O(\log_{2}(2u+1))$, and the number of interaction terms grows as $O((\log_{2}(2u+1))^{p})$ for the random interaction model, so the computational cost per sweep increases substantially. In particular, for $u=100$, as shown in Fig. 7(d), the heat bath method again shows slower energy convergence, and the optimal-transition Metropolis method shows an even clearer advantage over the other methods.
@@
Finally, we examine the results for the multilinear model. Note that, in this model, the heat bath method can determine the update of a variable in $O(1)$ time, and thus the difference in computational time among the transition probabilities is small. The log-encoded QUBO formulation is not plotted here, because the conversion process itself requires prohibitively long computation time. Figure 8(a) shows the results for $E^{\text{ML}}_{p=2,u=10^{6}}(\bm{z})$, where $u$ is extremely large. For short annealing times, the optimal-transition Metropolis method achieves the lowest energies, but it is overtaken by the heat bath method as the annealing time increases, and eventually, all methods converge to the same energy for sufficiently long annealing times. Figure 8(b) shows the results for a larger polynomial degree, $p=30,u=1$. As discussed at the end of Sect. 4.2, this model exhibits a large number of degenerate states with $E^{\text{ML}}_{p=30,u=1}(\bm{z})=0$, so convergence to the optimal solution requires a very long annealing time. Indeed, in Fig. 8(b), the energy does not reach the optimal value of 100-100, indicating that the annealing time is insufficient. In such short-time, the optimal-transition Metropolis method achieves the lowest energies among all transition probabilities. This tendency is also seen in Fig. 4(b).
In summary, directly solving the QUIO and HUIO formulations is more effective than converting them into QUBO and HUBO by log encoding. The heat bath method becomes slow when the variable bounds are large and high-order terms such as $z^{p}$ are present. Under these conditions, the optimal-transition Metropolis method shows better performance than the other methods, particularly for short annealing times.

### v2

@@ 4 Results
In this section, we present numerical results for SA with the transition probabilities described in Sect. 3.3. Our proposed algorithms are implemented in the open-source library OpenJij[openjij], and are freely available for use. All numerical experiments presented below were performed using OpenJij on a machine equipped with an Apple M1 Max processor and 64 GB of memory. In addition, we used the libraries JijModeling[cite_jijmodeling] and OMMX[cite_ommx_1, cite_ommx_2] to implement the QUIO and HUIO problems. Both libraries provide Python front ends with Rust back ends, enabling fast construction of problem instances.
@@
Let us summarize the basic properties of these models. The fully connected model $E^{\text{FC}}_{p,u}(\bm{z})$ consists of fully connected quadratic terms and on-site terms of degree $p$. Its ground-state energy is N(N+1)/2-N(N+1)/2. The multilinear model $E^{\text{ML}}_{p,u}(\bm{z})$ contains $p$-body multilinear interactions with periodic boundary conditions, $z_{i+N}=z_{i}$. Its ground-state energy is N-N. Both of these models have analytically known ground-state energies. In contrast, the ground-state energy of the random interaction model $E^{\text{RI}}_{p,u}(\bm{z})$ is generally difficult to determine, since it consists of interaction terms generated by selecting $p$ variable indices with repetition allowed, and assigning each term a coefficient $J_{i_{1},\ldots,i_{p}}$ sampled uniformly from [1,1][-1,1]. We set the number of interaction terms to 1000 so that the model becomes sparse. Note that this construction naturally produces on-site terms such as $z_{i}^{k}$ for some positive integer $k$.
@@
4.1 Annealing time under fixed sweeps
In this subsection we focus exclusively on the annealing time of SA for a fixed number of sweeps and we investigate how the annealing time, defined as the elapsed time from the start to the end of a single SA process, scales with the problem characteristics. In the following numerical experiments, SA was performed with 100 samples and 1000 sweeps, and the mean annealing time per single SA run was computed.
Let us first investigate how the variable range, $2u+1$, influences the annealing time, focusing on the fully connected model [Eq. (32)]. Figure 1 shows the annealing time as a function of $2u+1$, for $p=1$ and $p=2$. For $p=1$, the annealing time does not increase with $2u+1$ for any of the transition probabilities [Fig. 1(a)], because determining the update of a single variable does not depend on the variable range for all transition methods. For $p=2$, only the heat bath method requires $O(2u+1)$ time to determine the update of a single variable, because all $2u+1$ candidate values must be examined when performing the update. Consequently, as shown in Fig. 1(b), only the heat bath method exhibits annealing time that increases linearly with $2u+1$.
Figure 1: Annealing time as a function of variable range $2u+1$ for (a) $E^{\text{FC}}_{p=1,u}(\bm{z})$ and (b) $E^{\text{FC}}_{p=2,u}(\bm{z})$. Each data point represents the mean annealing time over 100 samples, with 1000 sweeps and $N=100$ variables. Error bars indicate the standard error.
Figure 2: Annealing time as a function of $p$ for (a) $E^{\text{FC}}_{p,u=1}(\bm{z})$ and (b) $E^{\text{ML}}_{p,u=10}(\bm{z})$. Each data point represents the mean annealing time over 100 samples, with 1000 sweeps and $N=100$ variables. Error bars indicate the standard error but are too small to be visible.
@@ 4.2 Energy as a function of temperature
Figure 3: Temperature $T$ dependence of the mean energy for (a) $E^{\text{FC}}_{p=2,u=1}(\bm{z})$, (b) $E^{\text{FC}}_{p=2,u=100}(\bm{z})$, (c) $E^{\text{RI}}_{p=4,u=1}(\bm{z})$, and (d) $E^{\text{RI}}_{p=4,u=100}(\bm{z})$. Each data point represents the mean energy at temperature $T$ during SA, averaged over 1000 samples, and the number of sweeps is 1000. The number of variables is $N=100$, and for $E^{\text{RI}}_{p,u}(\bm{z})$, the number of interactions is 1000. The black dotted line for (a) and (b) indicates the ground-state energy, 5050-5050.
@@
Similarly, Figs. 3(c) and 3(d) show the results for the random interaction model $E^{\text{RI}}_{p=4,u}(\bm{z})$. This model is a HUIO problem with fourth-order interactions and random coupling coefficients, and its characteristics therefore differ from those of the fully connected model $E^{\text{FC}}_{p=2,u}(\bm{z})$. Nevertheless, the results exhibit similar tendencies to those of $E^{\text{FC}}_{p=2,u}(\bm{z})$. The initial temperature appears appropriate, whereas the final temperature is also somewhat too low for this model as well.
@@
As for the differences among transition probabilities, the energy tends to decrease more readily when using the heat bath method than when using the Metropolis method [Figs. 3(b) and 3(d)], since it takes into account all possible variable updates instead of randomly selecting a single candidate. For the optimal-transition Metropolis method, the energy occasionally drops sharply when an optimal transition actually occurs, and thus its energy is generally lower than that of the other methods at the same temperature. In Fig.3(d), however, this method becomes trapped in a local minimum at low temperatures, and its final energy ends up slightly higher than that of the heat-bath method.
@@
Figure 4: Temperature $T$ dependence of the mean energy for $E^{\text{ML}}_{p=30,u=1}(\bm{z})$. Each data point represents the mean energy at temperature $T$ during SA, averaged over 1000 samples, and the number of sweeps is (a) 1000 and (b) 100000. The number of variables is $N=100$, and the black dotted line indicates the ground-state energy, 100-100.
@@
Next, we examine the results for the multilinear model $E^{\text{ML}}_{p=30,u=1}(\bm{z})$. For the case of 1000 sweeps [Fig. 4(a)], the energy is still decreasing even in the low-temperature region, suggesting that the final temperature is too high. In contrast, for 100000 sweeps [Fig. 4(b)], a region appears at low temperatures where the energy no longer decreases, indicating that the final temperature is sufficiently low. These results show that while the initial temperature is sufficiently high, the final temperature is not always appropriate and depends on the number of sweeps. In general, when the number of sweeps is small, the final temperature needs to be set lower. With fewer sweeps, variable updates are limited, so allocating more sweeps to the low-temperature region becomes important.
@@
We note that the multilinear model has a rather special nature. In each term of $E^{\text{ML}}_{p,u}(\bm{z})$, if any variable takes the value zero, the term contributes nothing to the energy. Because each term is a product of $p$ variables, the probability that a term is nonzero becomes extremely small for large $p$. For instance, when $u=1$, only $2^{p}$ configurations yield a nonzero value for a single term, whereas $3^{p}-2^{p}$ configurations yield zero. For $p=30$, this difference is enormous. As a result, there exist a vast number of configurations for which many multilinear terms evaluate to zero, and escaping from them requires a very large number of sweeps. Indeed, even with 100000 sweeps, the energy does not reach the ground-state value 100-100 [Fig. 4(b)], indicating that a much larger number of sweeps is required.
@@
In summary, the temperature-setting procedure in Sect. 3.4 is reasonable for the models studied here, although both the initial and final temperatures tend to be slightly conservative and are not necessarily optimal. We note also that the final temperature can be too high when the number of sweeps is small.
@@ 4.3 Energy as a function of sweeps
In this subsection, we examine how the energy decreases as a function of the number of sweeps. Since the computational cost per sweep depends on the choice of transition probability, the energy as a function of sweeps does not by itself quantify performance, but it helps reveal the characteristic behavior of each transition probability. In the following experiments, SA was performed with 1000 samples, and the mean energy for each fixed number of sweeps was computed.
We first consider the random interaction model [Eq. (34)] for $p=2$. When the variable range is small ($u=1$), the three transition probability methods exhibit almost no difference in performance [Fig. 5(a)]. In contrast, when the range is increased to $u=100$, the Metropolis method yields noticeably higher energies, particularly at small sweep regimes [Fig. 5(b)]. This behavior is consistent with the discussion in Sect. 3.3.1, which states that the Metropolis method performs poorly when the variable range is wide. It is also observed that the heat-bath method achieves slightly lower energies than the optimal-transition Metropolis method in this regime [Fig. 5(b)].
@@
The results for the multilinear model [Eq. (33)] show a somewhat different trend. For the extremely large variable range ($u=10^{6}$), the Metropolis method again performs poorly, but the heat-bath method exhibits a temporary degradation at small sweeps [Fig. 6(a)]. For the case of large polynomial degree ($p=30$), the system does not reach the ground state even after $10^{6}$ sweeps for any transition probability, and in this regime the optimal-transition Metropolis method yields the lowest energies [Fig. 6(b)].
Figure 6: Sweep dependence of the mean energy for (a) $E^{\text{ML}}_{p=2,u=10^{6}}(\bm{z})$ and (b) $E^{\text{ML}}_{p=30,u=1}(\bm{z})$. Each data point represents the mean energy averaged over 1000 samples. Error bars indicate the standard error but are too small to be visible. The number of variables is $N=100$ and the black dotted line indicates the ground-state energy, 100-100.
In summary, when compared at the same number of sweeps, the Metropolis method shows poor performance when the variable range is wide. For the other two methods, depending on the number of sweeps, either the heat-bath method or the optimal-transition Metropolis method can perform slightly better, but the difference between them remains small.
@@ 4.4 Energy as a function of annealing time
In this subsection, we examine the dependence of the optimized energy on the annealing time for different transition probabilities. This analysis aims to evaluate the performance of each transition probability, where better performance corresponds to achieving a lower energy within a shorter annealing time. In addition, we compare these results with those obtained by applying SA with the Metropolis method to QUBO and HUBO formulations that are constructed by converting the integer variables into binary ones using log encoding:
@@
Here, $m_{k}=\lceil\log_{2}(u_{k}-l_{k}+1)\rceil$ is the number of binary variables $b_{k,i}\in\{0,1\}$ corresponding to an integer variable $z_{k}$. Although the total computational cost should include the time required for binary encoding, we exclude it here to focus on the dependence of the achieved energy on the annealing time. It should be noted, however, that for models with higher-order interactions or large variable range, such binary conversion can require a substantial amount of time. In the following analysis, SA was performed with 1000 samples for several fixed numbers of sweeps, and the mean energy obtained in each case was plotted against the mean annealing time.
@@
First, we examine the results for the random interaction model [Eq. (34)]. Figure 7(a) shows the results for $p=2$ and $u=1$. When this model is converted into a QUBO formulation using the log encoding, the number of variables becomes twice as large and the number of interactions becomes four times larger, which increases the computational cost per sweep. As a result, SA applied to the log-encoded QUBO formulation yields higher energies at short annealing times, indicating poorer solution quality. In contrast, little difference is observed among the other transition probabilities, because when $u=1$, each variable can take only three states (1,0-1,0, and $1$), so the influence of the choice of transition probability becomes less significant. As the annealing time increases, the optimized energies obtained by all methods converge to similar levels, suggesting that all methods tend to reach the global optimum when sufficient annealing time is allowed.
Next, we examine the results for larger $u=100$ [Fig. 7(b)]. In this case, the difference between the log-encoded QUBO formulation and the directly solved QUIO model becomes larger than in the case of $u=1$. This is simply because both the number of variables and the number of interaction terms become much larger than in the case of $u=1$. The heat bath method also exhibits slower energy convergence with respect to annealing time, because approximately 1% of the interaction terms are squared terms such as $z^{2}$, and determining the updates for these variables requires $O(2u+1)$ time in the heat bath method. The optimal-transition Metropolis method shows lower energies in the short-time region and reaches the same final energies as the other methods in the long-time region, indicating that it provides the best overall performance. This is due to the effect of the optimal transitions defined in Eq. (25).
Similar tendencies are also observed for the HUIO case. As in the $p=2$ case, a clear difference is observed between the log-encoded HUBO formulation and the directly solved HUIO model [Figs. 7(c) and 7(d)]. In general, when the model is log-encoded, the number of variables increases by a factor of $O(\log_{2}(2u+1))$, and each original $p$-body interaction term expands into $O((\log_{2}(2u+1))^{p})$ binary terms. As a result, the computational cost per sweep increases substantially. In particular, for $u=100$, as shown in Fig. 7(d), the heat bath method again shows slower energy convergence, and the optimal-transition Metropolis method shows an even clearer advantage over the other methods.
@@
Finally, we examine the results for the multilinear model. Note that, in this model, the heat bath method can determine the update of a variable in $O(1)$ time, and thus the difference in computational time among the transition probabilities is small. The log-encoded QUBO formulation is not plotted here, because the conversion process itself requires prohibitively long computation time. Figure 8(a) shows the results for $E^{\text{ML}}_{p=2,u=10^{6}}(\bm{z})$, where $u$ is extremely large. For short annealing times, the optimal-transition Metropolis method achieves the lowest energies, but it is overtaken by the heat bath method as the annealing time increases, and eventually, all methods converge to the same energy for sufficiently long annealing times. Figure 8(b) shows the results for a larger polynomial degree, $p=30,u=1$. As discussed at the end of Sect. 4.2, this model contains many configurations for which multilinear terms take zero, so convergence to the optimal solution requires a very long annealing time. Indeed, in Fig. 8(b), the energy does not reach the optimal value of 100-100, indicating that the annealing time is insufficient. In this short-time regime, the optimal-transition Metropolis method achieves the lowest energies among all transition probabilities. This tendency is also seen in Fig. 4(b).
In summary, directly solving the QUIO and HUIO is more effective than converting them into QUBO and HUBO by log encoding. The heat bath method becomes slow when the variable range is wide and high-order terms such as $z^{p}$ are present. Under these conditions, the optimal-transition Metropolis method shows better performance than the other transition probabilities, particularly for short annealing times.